OpenStudy (anonymous):

Can someone tell me what I'm doing wrong? There were 200 tickets sold for a college basketball game. Tickets were $4.50 for students and $9 for adults. The total amount collect was $1485. How many of each type of ticket were sold?

OpenStudy (anonymous):

900a + 450s = 148500 -900 (a + s) = (200) -900 900a + 450s = 148500 -900a - 900s = -700 -450s = 147800 s = 328.44 (Completely incorrect)

OpenStudy (phi):

you are multiplying at this step -900 (a + s) = (200) -900 the right side is 200* -900 (not 200-900) you start with a+s=200 and multiply both sides by -900

OpenStudy (anonymous):

I also tried not including the .00 on every number 9a + 4.50s = 1485 -9a - 9s = -700 -4.5s = 785 s = 174.44 (also incorrect)

OpenStudy (anonymous):

Oh... I see... goddamn I hate myself xD

OpenStudy (phi):

Yes, I would stick with the smaller numbers, multiplying everything by 100 does not make it easier. In fact, I would divide everything by 9 in the 1st equation to get \[ \frac{9a}{9}+ \frac{4.5 s}{9}= \frac{1485}{9} \] which simplifies to a + \frac{1}{2}s = 165 and the problem is \[ a+s=200 \] \[a+\frac{1}{2}s= 165\] now these look easy to solve.

OpenStudy (phi):

When I make mistakes like this I try to remember what I was thinking and why I did what I did. It helps me to fix the "bug" in my thinking for next time.

OpenStudy (anonymous):

i think you should use simultaneous eqns.

OpenStudy (anonymous):

 x=number of student tickets y=number of adults tickets